OpenStudy (anonymous):

Fan and medal What number would you have to add to both sides to complete the square? x^2 + 12x = –6 A. –144 B. –36 C. 36 D. 144

OpenStudy (harsha19111999):

Can you elaborate the question?

OpenStudy (anonymous):

what do you mean?

OpenStudy (harsha19111999):

Expand the question

OpenStudy (anonymous):

Add a value to both sides of the equation that will be equal to the square of "half the coefficient" of x. So here you add (6)² to both sides: x² + 12x + 36 = -9 + 36 x² + 12x + 36 = 27 Now the perfect trinomial square can be factored into (x + 6)² and the problem becomes: (x + 6)² = 27 Take the square root of each side. You're setting up for solving for x: √((x + 6)²) = ±√(27) (x + 6) = ±√(27) (x + 6) = ±3√(3) Now for the positive part of the solution: (x + 6) = 3√(3) x + 6 = 3√(3) x = -6 + 3√(3) Then the negative: (x + 6) = -3√(3) x + 6 = -3√(3) x = -6 - 3√(3) The complete solution: x = -6 + 3√(3), -6 - 3√(3)

OpenStudy (solomonzelman):

\[x^2+bx=c\]Then adding\[\frac{b}{2}\] would make the left side a perfect square.

OpenStudy (solomonzelman):

I mean \[(~\frac{b}{2}~)~^2\]

OpenStudy (anonymous):

@ikatouni this doesn't help me AT ALL

OpenStudy (anonymous):

solve the splution and that's ur answer

OpenStudy (anonymous):

whats a splution?